OpenStudy (anonymous):

please help Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (anonymous):

I am not so certain, but i think it's along the line of -3 and then add 2. I am not so sure :| soryr. I am also studying about this.

OpenStudy (anonymous):

parallel means slope is the same the slope of \(y=mx+b\) is \(m\) so the slope of \(y=x+7\) is \(1\)

OpenStudy (anonymous):

y = x + 8 y = x - 4 y = x + 4 y = x - 8 those are the options they give me

OpenStudy (anonymous):

general equationn of a line passing through (x1,y1) is y-y1=m(x-x1). where m is the slope of the line.parallel line have the same slope,therefore the slope m is 1(as pointed out by satellite). x1,y1 are 6,-2 respectively.substitue these values in the general equation to get the answer

OpenStudy (anonymous):

so it 4

OpenStudy (anonymous):

no.take care of the signs and u get the answer

OpenStudy (anonymous):

y-6=1(x-(-2)) y-6=x+2 y=x+8